OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3? x + 3y = 10 x + 3y = –10 3x + y = 10 3x + y = –10

OpenStudy (anonymous):

you equation of slop

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

m=-3 x1=4 y1=-2

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

y+2=-3(x-4)

OpenStudy (anonymous):

y+2=-3x+12

OpenStudy (anonymous):

y+3x=10